Electric room heater draw current of 24A from 120V supply line what current will this room heater draw when connected to 240V supply line
1
Expert's answer
2019-08-12T10:05:17-0400

Let's first find the resistance of the heater from the Ohm's Law:


"R=\dfrac{V}{I} = \dfrac{120 V}{24 A} = 5 \Omega."

Then, using the Ohm's Law, we can find the current through the heater when it connected to 240 V supply line:


"I = \dfrac{V}{R} = \dfrac{240 V}{5 \Omega} = 48 A."

Answer:

"I = 48 A."